A company sells Gizmos to consumers at a price of $84 per unit. The cost to produce Gizmos is $22 per unit. The company will sell 10,000 Gizmos to consumers each year. The fixed costs incurred each year will be $200,000. There is an initial investment to produce the goods of $2,500,000 which will be depreciated straight line over the 6 year life of the investment to a salvage value of $0. The opportunity cost of capital is 12% and the tax rate is 39%. What is operating cash flow each year?
Answer #1

840,000 =10000*84 =10000*22+200000 =2500000/6 Sales Less: Costs Depreciation EBIT Less: Tax @ 32% Net income Add: Depreciatio

*Please rate thumbs up